# Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications

### v2

@@ 5.2 Distribution of Successful Perturbations
(i) QNLI Retrain Human
Figure 3: Plots of successful perturbations for all datasets when using TextFooler, showing the distribution of the number of instances across confidence bins.
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence# of Instances
(a) MRPC Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence
(b) MRPC Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence
(c) MRPC Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence# of Instances
(d) IMDB Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence
(e) IDMB Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence
(f) IMDB Retrain Human
(g) QNLI Original
(h) QNLI Retrain LLM
(i) QNLI Retrain Human
Figure 4: Plots of successful perturbations for all datasets when using DeepWordBug, showing the distribution of the number of instances across confidence bins.
6 Discussion
Our results demonstrate that accuracy remains consistent regardless of whether the additional synthetic samples are generated by a human or LLM, while the perturbation success rate exhibits varying degrees of change. The LLM-based approach has a greater capacity for reducing UUs compared to human-generated samples. However, it is important to note that human-generated samples have the potential to be more effective in certain circumstances. Overall, our findings highlight the strengths and limitations of both LLM- and human-based approaches, and their potential synergies in improving model performance and robustness.
6.1 LLM Capabilities
In our investigation, we found that the quality of responses obtained from human participants surpasses those of LLM in a limited number of cases, particularly when higher-order thinking skills were necessary. For example, in one of our perturbed samples, we had the date “June 15” in a sample changed to “John 15”. A high-performing crowdworker was able to discern the relation to the Bible verse that was most likely the cause of the UU, while the LLM was not. While the number of such high-quality responses is small, when present, their impact on reducing UUs can be significant and compensates for the many low-quality responses. Thus, we posit that the quality ceiling for humans is higher, although LLMs more consistently provide acceptable hypotheses and samples. It is worth noting that the leading reasons for poor-quality responses are either insufficient understanding of the task or a lack of motivation among participants Gadiraju et al. (2017).
6.2 Scalability and Ease of Use
Turning to the practical aspects of our study, we obtain important insights regarding the costs and time required for conducting human and LLM generalization experiments. The human study, involving 168 participants, incurred a cost of \$1072 at an hourly rate of \$12, whereas the LLM experiment had a significantly lower cost of \$46 for the same number of generalizations and samples generated. Though difficult to quantify, the collection of the data from humans via surveys also takes significantly longer than collecting LLM data. This reinforces the fact that the LLM-based approach is significantly more scalable, as it is not only cheaper but also yields data almost instantaneously, while obtaining data from humans comes with a significant delay. Even though the LLM-based procedure is clearly the most scalable, there might be some high-stakes or specialist applications where a human-based or hybrid procedure is most useful. These findings shed light on the resource implications of conducting human and LLM characterization studies, thereby aiding researchers in planning and allocating resources effectively.
7 Conclusion
We propose an approach to identify and mitigate blind spots in NLP models through generalization by humans or LLMs, followed by synthetic sample generation to address UUs and thus enhance model robustness. Our evaluation shows that our approach is successful in achieving a significant reduction in UUs across datasets while maintaining accuracy. We find that LLMs outperform humans in characterizing blind spots and that there is a significant task dependency to the quality of human contributions.
8 Limitations
It should be noted that the chosen method of perturbation inevitably influences the hypotheses generated by human participants and LLMs alike. Since our chosen perturbation approaches aim to swap in synonyms for vulnerable words or modify characters, found blind spots will only be due to these particular perturbations, thus biasing the characterization and generation in favor of such errors. Another limitation is the potential fallibility of humans or LLMs. Leveraging humans for the generalization and generation tasks risks ignoring blind spots that are not identifiable by humans. The same can be said for LLMs, which might miss human-identifiable blind spots. Further, contributions made by humans may be biased themselves due to personal biases or motivations. We try to address this through the previously discussed verification step. Finally, one notable challenge we encountered was the difficulty of definitively proving the origin of the responses provided by a small minority of participants, as their responses appear to resemble those generated by ChatGPT when asked to perform the same task. This echos the findings by other researchers, that crowdworkers are increasingly using LLMs to generate their responses Veselovsky et al. (2023). While this is just an observation that we cannot definitively prove, it raises the concern that some participants may not have authored their responses themselves, undermining the integrity of our study.
@@ Appendix A Synthetic Blind Spots
We use the synthetic blind spot study akin to a sanity check for our approach. As such, compared to the full natural blind spot study, we use a only a single task, a simpler model architecture, and make other simplifications to our mitigation process. We select an LSTM Hochreiter and Schmidhuber (1997) as our model of choice due to the absence of pretraining and apply the TextFooler perturbation method on the SA task. The LSTM used is the standard version of the Bi-LSTM provided by Morris et al. (2020).
@@
Table 6: Results of synthetic blind spot study for accuracy, perturbation success rate, and number of UUs before and after retraining for all LSTM model variants. The used perturbation method is TextFooler and the dataset is IMDB.
Blindspot Creation To assess whether our method can tackle existing synthetic blind spots we perform a type of Controlled Synthetic Data Check Nauta et al. (2023). We create synthetic blind spots by systematically excluding some data from training that have commonalities, namely containing a positive or negative term according to lexica by Liu et al. (2005). Here, we randomly subsample 600 of each as our selection of positive and negative terms, due to the extensive nature of the lexica.
@@
Blind Spot Mitigation After creating the synthetic blind spots through biasing, the authors perform the generalization procedure and provide handcrafted hypotheses that precisely describe these, similar to golden labels. To generate the new samples from our handcrafted hypotheses, we prompt ChatGPT to generate movie review-related sentences (to fit the chosen task) that follow a given hypothesis. This was done in an attempt to simplify the procedure by taking advantage of human strengths, generalization and extrapolative thinking, and LLM strengths, low-cost text generation, simultaneously.
A.1 Synthetic Blind Spot Study Results
The mitigation results of this human-LLM approach for our Controlled Synthetic Data Check can be seen in table 6.
@@ A.1 Synthetic Blind Spot Study Results
Outcomes of Biasing As can be seen in the first column of table 6, before retraining, the overall test accuracy declines in line with the degree to which the train set is biased. Interestingly, the percentage of successful perturbations by TextFooler, i.e., the percentage of successful label flips, closely follows the overall accuracy. This mirrors the findings of Tsipras et al. (2019), that there is a strong relationship between high accuracy and brittleness – or a lack of robustness. The number of occurring UUs as a result of the perturbation does not follow this trend, instead increasing as the training data becomes more biased, as expected. This poses an interesting optimization problem since the model becomes most robust in general terms, i.e., the successful perturbation percentage falls, but simultaneously there is a significant uptick in blind spots as the training sets become more biased.
@@
Blind Spot Mitigation The effect of retraining on the overall accuracy and perturbation success rate is minimal, with accuracy changing by no more than $\pm$ 1% and perturbation success rate changing no more than $\pm$ 2%. However, the number of found UUs decreases drastically due to the retraining, with reductions of 73.93%, 80.77%, and 74.75% for the biased P, N, and PN models, respectively. The clean and randomly biased models also show a reduction, though less significant at 54.55% and 31.49%, respectively. These results confirm that our method can be used to target synthetic blind spots found in biased models through the use of hypotheses and generated instances, without significantly affecting the performance or general robustness of the model.
Appendix B Perturbation Statistics
To add additional context to the perturbation performed, we supply the detailed attack statistics across all performed perturbations. Specifically, we report Original Accuracy and Accuracy Under Attack are reported, which are the classifier accuracy on its own and while under attack. Further, Attack Success Rate is shown, which is the percentage of successful perturbation attempts to failed ones. Finally, we report the number of Perturbed Words, the percentage of words that are perturbed, the Words per Input, the average number of words per input, and the Average Number of Queries, which is how many tries it took the perturbation method to find the best attack. The attack statistics for TextFooler attacks are shown in table 7 while the ones for DeepWordBug attacks are shown in table 8.
[…]
Original Accuracy (%) 82.38 81.57 81.58 94.84 95.40 94.43 89.88 89.31 89.42
Accuracy Under Attack (%) 9.80 17.40 12.99 10.18 10.44 19.21 8.91 11.67 14.89
Attack Success Rate (%) 71.83 64.87 68.29 88.46 93.18 63.85 87.35 86.80 78.84
Perturbed Words (%) 7.70 9.9 8.51 4.59 7.62 9.02 6.12 8.80 9.57
Words per Input 39.3 39.3 39.3 230.0 230.0 230.0 37.9 37.9 37.9
Avg. Number of Queries 51.40 68.62 55.17 185.24 184.94 198.31 49.38 51.27 56.11
Table 7: Perturbation statistics across datasets and models for attacks with TextFooler. Subscripts O, L, and H denote the original, LLM-retrained, and human-retrained models, respectively.
[…]
Original Accuracy (%) 82.38 82.23 82.10 95.40 95.41 95.74 89.88 89.38 89.38
Accuracy Under Attack (%) 7.78 13.73 11.94 9.54 21.43 15.32 8.21 9.90 7.30
Attack Success Rate (%) 72.00 70.38 72.64 59.41 50.59 79.70 77.54 79.74 82.08
Perturbed Words (%) 8.47 9.18 9.03 6.43 8.11 13.09 7.99 8.32 11.03
Words per Input 39.3 39.3 39.3 230.0 230.0 230.0 37.9 37.9 37.9
Avg. Number of Queries 56.92 64.37 58.61 199.32 211.65 201.44 34.91 33.53 49.09
Table 8: Perturbation statistics across datasets and models for attacks with DeepWordBug. Subscripts O, L, and H denote the original, LLM-retrained, and human-retrained models, respectively.
Appendix C User Study Structure
@@
We use Prolific as a crowdsourcing platform for all our participants. Below, we present the structure followed by all survey participants for the generalization user study, consisting of an initial disclaimer, an instruction set, examples, and finally the questions. Here, we use the abstraction and exploration assignments on the IMDB dataset as an example. The workflow is very similar between the different generalization assignments and datasets (MRPC, IMDB, or QNLI), with only slight differences in the wording between the surveys to fit the task and dataset used, as they all present the crowd worker with some input and result in plain text output. For the generation assignment, crowdworkers are asked to perform the same steps, with relevant examples related to the structure of the dataset being shown, before finally contributing usable samples based on shown hypotheses.
C.1 Abstraction on IMDB
@@
  • “Welcome to the Hypothesis Exploration Survey! Please carefully read the following: You are invited to participate in our research study. This study is fully sanctioned by our governing ethics body, as is the handling and storing of the resulting data. This research study aims to use your creativity and generalization ability to come up with new abstractions. It will take you approximately 25 minutes to complete. As with any online activity, the risk of a breach is always possible. To the best of our ability, your answers in this study will remain confidential. We will minimize any risks by making this survey completely anonymous. Therefore, please do not provide any personal information anywhere. The anonymous results might be shared publicly in the future. Participation in this study is entirely voluntary, and you can withdraw anytime. Feel free to contact us with any questions or feedback you might have.”
@@
    The pattern that caused the wrong prediction may be related to the substitution of the word ""story"" with its synonym ""narrative"".
@@
    Several words have been misspelled in the samples, all related to the letter ""o"". Either more letters are added ""oo"" or the letter is substituted with a number ""0"" that looks similar, making it easy to misread.
@@
    O: <original sentence> P: <perturbed sentence>
@@
Appendix D Used LLM Prompts
@@ D.1 Abstraction Prompt
    Hypothesis: <<hypothesis>>
    Reasoning: <<reasoning>>”
D.2 Exploration Prompt
@@
    Hypothesis: <<hypothesis>>
    Reasoning: <<reasoning>>”
D.3 Generation Prompt
@@
    The new sample should be varied and detailed. Follow the logic laid out in the given hypothesis and follow the format of the sample pair (O and P) exactly. Also include whether the new sample should be given a <<positive>> or <<negative>> label for the task: {task}.
@@
    Sample: <<sample>>
    Label: <<label>>
    Reasoning: <<reasoning>>”

### v3

@@ 4.2.2. Observation 6: Obtaining samples via LM is easier and more cost effective
(d) Relabel Baseline
Figure 4. Plots of prediction confidence per misclassified sample for BERT on QNLI dataset when using TF as a perturbation technique, showing the distribution across confidence bins. The distribution of the prediction confidences is altered by the retraining, regardless of how it was performed. Our method is able to lower the number of high-confidence classifications, especially those at the highest of confidences, improving model calibration.
5. Related Work
In this section, we briefly review relevant prior research on approaches to high confidence misclassifications, as well as how others have tried to avoid such model behaviour.
5.1. Unknown Unknowns
Attenberg et al. (2015) introduce the concept of querying humans to find UUs in a game-like setting and show that there were patterns to the found UUs. Vandenhof (2019) proposes an approach to identify UUs where human-interpretable decision rules are learned to approximate how a model makes high-confidence predictions. Crowdworkers then contradict these rules by finding an instance that would classify as a UU. Cabrera et al. (2021) explore the use of crowdworkers to generate failure reports for computer vision models to describe how or why the model failed. Han et al. (2021) propose an approach where crowdworkers continuously extend a dataset with relabeled UUs, on which the chosen model is iteratively trained. Instead, we go beyond simple relabeling and characterize found blind spots and explore new, previously unseen blind spots. There are also algorithmic approaches to finding UUs, such as Lakkaraju et al. (2017), who propose utilizing an explore-exploit approach to find groups of UUs. Bansal and Weld (2018) extend this by proposing a utility model that rewards the degree to which the found UUs cover a sample distribution, thus encouraging the discovery of new blind spots. Instead, we do not find the UUs algorithmically, but instead use an LM or crowdworkers to find existing UUs, extrapolate from these to unseen UUs, and generate synthetic data targeting both of these.
5.2. Model Calibration and Robust Training
The concept of UUs and blind spots is connected to model calibration (Guo et al., 2017; Minderer et al., 2021; Tian et al., 2023). A model that is well-calibrated will have its prediction confidence aligned with the likelihood of the correctness of the prediction and, as such, a model with blind spots is a poorly calibrated model. In the case where the UUs are specifically generated through adversarial attacks, illumination of model blind spots is also related to robust training. UUs that populate these blind spots, when created by such attacks, may be identified as adversarial examples (Ribeiro et al., 2018; Wallace et al., 2019; Wang et al., 2020). This underscores the relationship between our proposed method and robust training practices with the aim of improving the robustness of the model (Madry et al., 2018; Pang et al., 2021). Our method focuses not on general robustness but rather on high-confidence misclassifications and is not limited to just adversarial samples, as we consider UUs that occur naturally without perturbation as well.
Several approaches have been proposed to utilize synthetic data to expand training sets (Puri et al., 2020; Claveau et al., 2021).  He et al. (2022) explore few-shot prompting LMs to generate task specific synthetic training data. Unlike prior work, we propose a method to generate targeted synthetic data with the purpose of eliminating blind spots that lead to high confidence misclassifications.
6. Conclusion
We propose a method to identify and mitigate blind spots in classification models by leveraging human- and LLM-generated generalizations, followed by synthetic sample generation to target UUs and enhance model robustness. Our evaluation demonstrates that our method is effective at addressing model blind spots and achieves a significant reduction in UUs across datasets, while not altering the general performance of the model and therefore maintaining accuracy. Our study sheds light on the notable task dependency of the human ability to characterize blind spots and generate new data and how this ability compares to that of an LM. Future work will focus on optimizing the balance between accuracy and robustness to further enhance model performance.
@@ Appendix A Synthetic Blind Spots
We use the synthetic blind spot study akin to a sanity check for our approach. As such, compared to the full natural blind spot study, we use a only a single task, a simpler model architecture, and make other simplifications to our mitigation process. We select an LSTM (Hochreiter and Schmidhuber, 1997) as our model of choice due to the absence of pretraining and apply the TF perturbation method on the SA task. The LSTM used is the standard version of the Bi-LSTM provided by Morris et al. (2020).
@@
Table 3. Results of synthetic blind spot study for accuracy, perturbation success rate, and number of UUs before and after retraining for all LSTM model variants. The used perturbation method is TF and the dataset is IMDB.
A.1. Blindspot Creation and Mitigation
To assess whether our method can tackle existing synthetic blind spots we perform a type of Controlled Synthetic Data Check (Nauta et al., 2023). We create synthetic blind spots by systematically excluding some data from training that have commonalities, namely containing a positive or negative term according to lexica by Liu et al. (2005). Here, we randomly subsample 600 of each as our selection of positive and negative terms, due to the extensive nature of the lexica.
@@ A.1. Blindspot Creation and Mitigation
After creating the synthetic blind spots through biasing, the authors perform the generalization procedure and provide handcrafted hypotheses that precisely describe these, similar to golden labels. To generate the new samples from our handcrafted hypotheses, we prompt ChatGPT to generate movie review-related sentences (to fit the chosen task) that follow a given hypothesis. This was done in an attempt to simplify the procedure by taking advantage of human strengths, generalization and extrapolative thinking, and LLM strengths, low-cost text generation, simultaneously.
A.2. Synthetic Blind Spot Study Results
@@
The mitigation results of this human-LLM approach for our Controlled Synthetic Data Check can be seen in table 3. As can be seen in the first column of table 3, before retraining, the overall test accuracy declines in line with the degree to which the train set is biased. Interestingly, the percentage of successful perturbations by TF, i.e., the percentage of successful label flips, closely follows the overall accuracy. This mirrors the findings of Tsipras et al. (2019), that there is a strong relationship between high accuracy and brittleness – or a lack of robustness. The number of occurring UUs as a result of the perturbation does not follow this trend, instead increasing as the training data becomes more biased, as expected. This poses an interesting optimization problem since the model becomes most robust in general terms, i.e., the successful perturbation percentage falls, but simultaneously there is a significant uptick in blind spots as the training sets become more biased.
@@
The effect of retraining on the overall accuracy and perturbation success rate is minimal, with accuracy changing by no more than $\pm$ 1% and perturbation success rate changing no more than $\pm$ 2%. However, the number of found UUs decreases drastically due to the retraining, with reductions of 73.93%, 80.77%, and 74.75% for the biased P, N, and PN models, respectively. The clean and randomly biased models also show a reduction, though less significant at 54.55% and 31.49%, respectively. These results confirm that our method can be used to target synthetic blind spots found in biased models through the use of hypotheses and generated instances, without significantly affecting the performance or general robustness of the model.
Appendix B User Study for Human Computation
@@
We use Prolific as a crowdsourcing platform for all our participants. Below, we present the structure followed by all survey participants for the generalization user study, consisting of an initial disclaimer, an instruction set, examples, and finally the questions. Here, we use the abstraction and extrapolation assignments on the IMDB dataset as an example. The workflow is very similar between the different generalization assignments and datasets (MRPC, IMDB, or QNLI), with only slight differences in the wording between the surveys to fit the task and dataset used, as they all present the crowd worker with some input and result in plain text output. For the generation assignment, crowdworkers are asked to perform the same steps, with relevant examples related to the structure of the dataset being shown, before finally contributing usable samples based on shown hypotheses.
B.1. Abstraction on IMDB
@@
  • “Welcome to the Hypothesis Extrapolation Survey! Please carefully read the following: You are invited to participate in our research study. This study is fully sanctioned by our governing ethics body, as is the handling and storing of the resulting data. This research study aims to use your creativity and generalization ability to come up with new abstractions. It will take you approximately 25 minutes to complete. As with any online activity, the risk of a breach is always possible. To the best of our ability, your answers in this study will remain confidential. We will minimize any risks by making this survey completely anonymous. Therefore, please do not provide any personal information anywhere. The anonymous results might be shared publicly in the future. Participation in this study is entirely voluntary, and you can withdraw anytime. Feel free to contact us with any questions or feedback you might have.”
@@
    The pattern that caused the wrong prediction may be related to the substitution of the word ””story”” with its synonym ””narrative””.
@@
    Several words have been misspelled in the samples, all related to the letter ””o””. Either more letters are added ””oo”” or the letter is substituted with a number ””0”” that looks similar, making it easy to misread.
@@
    O: {original sentence} P: {perturbed sentence}
@@
Appendix C Used LLM Prompts
@@ C.1. Abstraction Prompt
    Hypothesis: {hypothesis}
    Reasoning: {reasoning}”
C.2. Extrapolation Prompt
@@
    Hypothesis: {hypothesis}
    Reasoning: {reasoning}”
C.3. Generation Prompt
  • “There is a sentence pair, with one original sample (O) and a perturbed one (P), which is similar but had some things changed. These changes may be related to a pattern related to semantics, syntax, specific words, or something else that leads to them being the reason the sample is misclassified by a classification algorithm. This misclassification is made at a high level of confidence.
@@ C.3. Generation Prompt
    The new sample should be varied and detailed. Follow the logic laid out in the given hypothesis and follow the format of the sample pair (O and P) exactly. Also include whether the new sample should be given a (positive) or (negative) label for the task: {task}.
@@
    Sample: {sample}
    Label: {label}
    Reasoning: {reasoning}”
Appendix D Perturbation Statistics and Visualization
[…]
Original Accuracy (%) 82.38 81.57 81.58 82.49 94.84 95.40 94.43 93.94 89.88 89.31 89.42 88.24
Accuracy Under Attack (%) 9.80 17.40 12.99 10.42 10.18 10.44 19.21 10.22 8.91 11.67 14.89 9.97
Attack Success Rate (%) 71.83 64.87 68.29 69.65 88.46 93.18 63.85 85.34 87.35 86.80 78.84 84.92
Perturbed Words (%) 7.70 9.9 8.51 7.98 4.59 7.62 9.02 5.50 6.12 8.80 9.57 7.33
Words per Input 39.3 39.3 39.3 39.3 230.0 230.0 230.0 230.0 37.9 37.9 37.9 37.9
Avg. Number of Queries 51.40 68.62 55.17 57.86 185.24 184.94 198.31 186.37 49.38 51.27 56.11 53.27
Table 4. Perturbation statistics across datasets and models for attacks with TF using BERT. Subscripts O, L, H, R denote the original, LM-retrained, human-retrained, and relabeled models, respectively.
[…]
Original Accuracy (%) 82.38 82.23 82.10 82.55 95.40 95.41 95.74 94.26 89.88 89.38 89.38 88.98
Accuracy Under Attack (%) 7.78 13.73 11.94 10.42 9.54 21.43 15.32 12.51 8.21 9.90 7.30 8.67
Attack Success Rate (%) 72.00 70.38 72.64 72.35 59.41 50.59 79.70 56.87 77.54 79.74 82.08 79.27
Perturbed Words (%) 8.47 9.18 9.03 8.91 6.43 8.11 13.09 9.37 7.99 8.32 11.03 8.31
Words per Input 39.3 39.3 39.3 39.3 230.0 230.0 230.0 230.0 37.9 37.9 37.9 37.9
Avg. Number of Queries 56.92 64.37 58.61 58.23 199.32 211.65 201.44 204.12 34.91 33.53 49.09 35.75
Table 5. Perturbation statistics across datasets and models for attacks with DWB using BERT. Subscripts O, L, H, R denote the original, LM-retrained, human-retrained, and relabeled models, respectively.
[…]
Original Accuracy (%) 90.84 89.86 90.20 90.61 95.20 94.96 94.67 94.86 90.08 89.58 89.16 89.90
Accuracy Under Attack (%) 13.85 18.31 12.43 14.09 20.97 18.22 15.09 17.55 12.64 15.29 14.53 13.67
Attack Success Rate (%) 68.70 65.24 69.54 66.89 71.32 75.64 78.31 70.55 83.42 79.12 75.87 81.34
Perturbed Words (%) 9.23 8.12 9.68 8.97 6.45 7.54 10.88 8.36 7.34 8.69 9.11 7.92
Words per Input 39.3 39.3 39.3 39.3 230.0 230.0 230.0 230.0 37.9 37.9 37.9 37.9
Avg. Number of Queries 53.92 62.34 57.92 55.76 191.34 192.85 198.21 194.43 48.22 49.98 52.89 50.76
Table 6. Perturbation statistics across datasets and models for attacks with TF using Llama 2. Subscripts O, L, H, R denote the original, LM-retrained, human-retrained, and relabeled models, respectively.
[…]
Original Accuracy (%) 90.66 89.73 89.91 90.73 95.33 95.13 94.90 95.10 90.72 90.10 89.73 90.60
Accuracy Under Attack (%) 16.35 14.79 13.87 15.68 21.78 20.32 19.12 22.19 11.78 10.95 14.28 12.44
Attack Success Rate (%) 66.40 63.89 67.56 65.78 70.42 68.55 71.32 74.65 79.78 77.24 82.43 80.34
Perturbed Words (%) 9.11 8.76 9.02 8.86 7.18 6.92 11.54 9.29 8.06 9.11 10.24 8.76
Words per Input 39.3 39.3 39.3 39.3 230.0 230.0 230.0 230.0 37.9 37.9 37.9 37.9
Avg. Number of Queries 60.22 65.14 62.03 61.76 203.56 199.42 204.29 208.23 45.29 43.87 50.77 47.83
Table 7. Perturbation statistics across datasets and models for attacks with DWB using Llama 2. Subscripts O, L, H, R denote the original, LM-retrained, human-retrained, and relabeled models, respectively.
To add additional context to the perturbation performed, we supply the detailed attack statistics across all performed perturbations. Specifically, we report Original Accuracy and Accuracy Under Attack are reported, which are the classifier accuracy on its own and while under attack. Further, Attack Success Rate is shown, which is the percentage of successful perturbation attempts to failed ones. Finally, we report the number of Perturbed Words, the percentage of words that are perturbed, the Words per Input, the average number of words per input, and the Average Number of Queries, which is how many tries it took the perturbation method to find the best attack. For BERT, the attack statistics for TF attacks are shown in table 4 while the ones for DWB attacks are shown in table 5. For Llama 2 7B, the attack statistics for TF attacks are shown in table 6 and for DWB in table 7.